# An Exposition of Pathfinding Strategies Within Lightning Network Clients

### v1

@@ Abstract
The Lightning Network is a peer-to-peer network designed to address Bitcoin’s scalability challenges, facilitating rapid, cost-effective, and instantaneous transactions through bidirectional, blockchain-backed payment channels among network peers. Due to a source-based routing of payments, different pathfinding strategies are used in practice, trading off different objectives for each other such as payment reliability and routing fees. This paper explores differences within pathfinding strategies used by prominent Lightning Network node implementations, which include different underlying cost functions and different constraints, as well as different greedy algorithms of shortest path-type. Surprisingly, we observe that the pathfinding problems that most LN node implementations attempt to solve are NP-complete, and cannot be guaranteed to be optimally solved by the variants of Dijkstra’s algorithm currently deployed in production. Through comparative analysis and simulations, we evaluate efficacy of different pathfinding strategies across metrics such as success rate, fees, path length, and timelock. Our experiments indicate that the strategies used by LND tend to be advantageous in terms of payment reliability, Eclair tends to result in paths with low fees, and that LDK exhibits average reliability with larger fee levels for smaller payment amounts; furthermore, CLN stands out for its minimal timelock paths. Additionally, we investigate the impact of Lightning node connectivity levels on routing efficiency. The findings of our analysis provide insights towards future improvements of pathfinding strategies and algorithms used within the Lightning Network.
@@ I Introduction and Background
By leveraging the trust and security of the Bitcoin blockchain in its payment channels, the Lightning Network offers bitcoin-denominated fast, cost-effective, and scalable transactions. Within the LN, users establish bidirectional payment channels, which collectively form a peer-to-peer network. These channels operate as 2-of-2 multisignature contracts between two parties (nodes) that hold bitcoin. Through these channels, users, which serve as nodes within the LN, can conduct off-chain transactions. Settlement of these off-chain transactions on the Bitcoin blockchain occurs only when necessary, such as when a user opts to close a channel. This architecture renders the Lightning Network scalable and enables instant payments [5].
@@
Multi-hop payments are facilitated through source routing, wherein the source locates a path to the receiver for payment routing. Initially, the sender receives a cryptographic hash from the recipient. For the sender’s payment to succeed, the recipient, upon receiving the HTLC, must return the pre-image along the HTLC-routed path within a specified timelock. Consequently, upon payment initiation, a series of HTLCs are established along the intended path. Each HTLC represents a commitment of funds from the preceding node to the next node, contingent upon the fulfillment of cryptographic proof. Once all contracts are established, the receiver shares the pre-image of the hash with its predecessor hop, which forwards it along the path until it reaches the sender, thereby resolving HTLCs along the path [8]. LN nodes can route payments only if they have sufficient balance on their side of the channel. A key practical challenge for successful payments within the LN is that the sender node selecting a payment path only possesses information about the capacity of channels, but not about the balance distribution within them [9], which might lead to the selection of payment path that cannot successfully route the payment amount.
In order to discover paths for routing payments, the sender node utilizes a pathfinding algorithm. Popular and widely used pathfinding algorithms have been developed by various LN implementations, including Lightning Network Daemon (LND) [10], Core Lightning (CLN) [11], Lightning Development Kit (LDK) [12], and Eclair [13]. While mostly using a variant of a shortest-path algorithm, each of these node implementations differs in its pathfinding strategy in substantial details such as in the weight function that establishes underlying path costs. With their different choices, these LN client implementations adopt diverse approaches to the pathfinding process with different trade-offs regarding the fees incurred by payments, the typical length of payment paths, incurred delays and payment reliability. Only little research has been dedicated to selecting the appropriate client (or pathfinding strategy) based on different need profiles regarding these trade-offs inherent to payment delivery within LN.
@@ II Related Work
The study by [14] concentrates on the reliability and privacy of LN payments by introducing a mathematical framework to model uncertainty in channel balances using probability theory. Furthermore, [15] explores the privacy properties of the Lightning Network through simulations. [16] investigates the structure of the Lightning Network and the routing algorithms used by LND, CLN and Eclair. It then explores the effects of routing choices by analyzing the feasibility of Denial-of-Service attacks and route hijacking across these three implementations. In our work, we specifically analyze the effects of design elements in the pathfinding algorithm and evaluate their empirical performance from an operational perspective.
[17] examines the potential presence of the Braess paradox in off-chain networks and proposes solutions to mitigate its impact through analysis employing three routing algorithms, including LND and CLN. [18] analyzes the Lightning Network’s topology and introduces a strategy called ProfitPilot for connecting new nodes in payments by encouraging cycle creation in payment channel networks. Studies such as [19], [15], and [20] present topological analyses of the Lightning Network. In our study, we utilize insights into the Lightning Network topology, including the node connectivity levels, to analyze the performance of pathfinding strategies for different network participants.
@@
Beyond the scope of our exposition are payment delivery strategies based on (atomic) multi-part payments [6, Chapter 12], which have been introduced as an extension to the original LN protocol [24]. [25] argues that multipart payment delivery benefits from integrating payment probability and the formulation via minimum cost flows instead of following a shortest-path framework. [26] proposes a variation of Boomerang which requires less computing power and latency for faster multi-path payments. While multi-part payments are very promising, they do currently not (yet) represent the mainstream practice in LN, where single-path payments still dominate. Thus, our study focuses on single-path payments, but the conducted analysis could be extended to multi-part payments in future work.
@@ III Pathfinding and Route Selection
For the purpose of pathfinding, the Lightning Network can be modelled as a directed graph $G=(V,E)$, in which each network participant is represented as a node or vertex $v\in V$, and a payment channel between two nodes $u,v\in V$ is represented as the pair of edges $e=(u,v)\in E$ and $e=(v,u)\in E$. Each edge in the graph is associated with specific channel attributes such as fees, capacity, timelock delta, age, and information amount the maximal and minimal amount of currency units that can be settled in an HTLC. These channel attributes are also utilized as design elements for pathfinding purposes. It is important to note that the LN graph evolves dynamically over time–the number of nodes and channels changes through channel opening and closing transactions, and beyond that, node operators adapt channel parameters such as fee parameters and communicate those changes via a gossip protocol to other network participants.
@@
When the payment is initiated, the source node needs to find a path to the receiver node using pathfinding strategies on the LN graph. Existing LN clients employ instantiations of shortest path algorithms, such as Dijkstra’s algorithm, modified Dijkstra-style algorithms or Yen’s $K$-shortest path algorithm, to find the path between specified source and receiver nodes [27], [28], [29] and [30]. We refer to Section V for a discussion of these variants in the context of their use in LN implementations.
@@ IV Weight Function Design Elements
Given the a sender node $s\in V$ and a recipient node $r\in V$, shortest path algorithms aim to find a path $p=(e_{1},\ldots,e_{\operatorname{length}(p)})$ of edges $e_{i}\in E$, where $s$ is the initial node of $e_{1}$ and $r$ is the terminal node of $e_{\operatorname{length}(p)}$, which minimizes a total path cost $c(p)$ given as the sum of edge weights given by a function $\operatorname{weight}:E\to\operatorname{\mathbb{R}}_{\geq 0}$ that maps to non-negative real numbers $\operatorname{\mathbb{R}}_{\geq 0}$, such that
@@
cf. [31, Chapter 22]. Recalling that an edge in the LN graph corresponds to a payment channel, this means that current LN client implementations, which use the shortest path algorithms to find their payment paths (see Section V), implicitly assume additive path costs in the sense of (1).
@@ IV-A Fees
When routing payments from a source $s$ to a recipient node $r$, the payment path $p$ typically involves intermediary nodes and multiple hops. These intermediary nodes charge fees for each successfully settled transaction they participate in. These fees can be subdivided into two distinct parts: the base fee ($\mathrm{b}_{e}$), which remains constant regardless of the transaction size, and the proportional fee or fee rate ($\mathrm{f}_{e}$), which varies based on the amount being forwarded through the node. Consequently, the total fee charged by an intermediary node for forwarding a specific payment amount ($\mathrm{amt}_{e}$) is determined by the combination of these two components, such that
$\mathrm{fee}_{e}=\mathrm{b}_{e}+\mathrm{amt}_{e}\cdot f_{e}.$ (3)
Since it is reasonable for the source node to try to find a path with fees as low as possible, one of the weight terms $a_{e}^{i}$ in (2) can be chosen as $a_{e}^{i}=\mathrm{fee}_{e}$ defined above. This choice is a primary design consideration in all LN implementations when formulating their weight functions. Since fees are directional, i.e., they can vary depending on the direction of payment flow between the same two nodes $u,v\in V$, a modeling of the LN as a directed graph instead of an undirected one is more appropriate.
@@ IV-B Timelock
Once a specific number of blocks is reached — referred to as a timelock — each incoming HTLC in a channel expires. This timelock serves as a time buffer within which the node aims to complete all the HTLC exchanges for a given transaction. For the length of the timelock, funds are locked in the payment channels.
The total timelock of an HTLC, in direction from sender to recipient, depends on the timelock of the subsequent HTLC along a payment path plus the timelock difference value $\Delta^{\text{CLTV}}_{e}$, called timelock delta or CLTV expiry delta [^1] CLTV is an acronym of the Bitcoin script operator OP_CHECKLOCKTIMEVERIFY [6, Glossary]. that is set by the node $v$ at the outgoing side of the channel $e=(u,v)$ [6, Chapter 8]. In order to ensure that funds are locked for a shorter duration, it is reasonable for routing clients to prioritize the inclusion of channels with lower timelock deltas into the payment path compared to channels with higher timelock deltas.
@@ IV-C Success Probability
Due to the fact that payment paths in which the desired payment amount can be settled reliably are preferred, it is common to take an estimate of a success probability, which estimates the likelihood of a successful payment through a payment path, into account when searching for suitable payment paths $p$. In this subsection, we first discuss two different approaches for incorporating payment path success probabilities into a model for a total path cost, and then shed light on different approaches for modelling individual channel success probabilities.
@@ IV-C1 Path Success Probability
In existing literature on LN pathfinding, there are two different approaches for incorporating a path success probability estimate $P_{p}$ into a total path cost $c(p)$ of a payment path $p$.
The first, which is used in LND [32, 33] implies a success probability penalization that is proportional to the inverse $P_{p}^{-1}$ of the path success probability $P_{p}$. It is based on the following derivation.
@@
We note that in [32, 33], the derivation above is conducted assuming the existence of a virtual cost $c_{\text{attempt}}$ only in case of a payment failure, but both assumptions lead to the same objective (4). In (4), the fee cost term $F_{p}$ fits well into the additive framework of (1) due the fact that routing fees accumulate additively along the path such that $F_{p}=\sum_{e\in p}\operatorname{weight_{a}}(e)$, where $\operatorname{weight_{a}}:E\to\operatorname{\mathbb{R}}_{\geq 0}$ is a non-negative, additive weight function.
However, for second summand of (4), the situation is different as an additive decomposition of $\frac{c_{\text{attempt}}}{P_{p}}$ across edges can be mathematically not justified. On the other hand, a reasonable assumption is that a payment across a path $p$ is successful if and only if the routing across all channels $e\in p$ is successful, who themselves are assigned channel success probabilities $P_{e}$. If we assume that the routing success across one channel $e\in p$ is independent from the routing success across any other channels along each path, we have that
$\begin{split}P_{p}&=P\left(\{\text{payment across path }p\text{ is successful}\}\right)\\ &=P\left(\cap_{e\in p}\{\text{payment across channel }e\text{ is successful}\}\right)\\ &=\prod_{e\in p}P_{e},\end{split}$
@@
Unfortunately, it is not possible to use a standard shortest path algorithm such as Dijkstra’s algorithm to find the path with the lowest additive-plus-multiplicative cost (5) as these algorithms all assume additive cost as in (1) [31, Chapter 22].
The LND LN node implementation mitigates this issue by modifying the priority queue used in Dijkstra’s algorithm to be adapted to (5). However, as we discuss in Section V, it can be shown that such a modified Dijkstra is not guaranteed to find the optimal path with respect to (5).
@@
which leads to a fully additive cost of the type (1) motivates the choice of a weight term $a_{e}^{i}=-\log(P_{e})$ in (2). Since the edge success probabilities $P_{e}$ are between 00 and $1$, these weight terms are always non-negative. If used without any other weight terms (e.g., fee-related ones), the resulting problem type corresponds to finding a maximally reliable path and is known to be equivalent to a shortest path problem [34, Exercise 4.39]. CLN, LDK and Eclair use this logarithmic success probability term in their weight function to prioritize paths with higher success probability within their pathfinding algorithms.

### v2

@@ Abstract
The Lightning Network is a peer-to-peer network designed to address Bitcoin’s scalability challenges, facilitating rapid, cost-effective, and instantaneous transactions through bidirectional, blockchain-backed payment channels among network peers. Due to a source-based routing of payments, different pathfinding strategies are used in practice, trading off different objectives for each other such as payment reliability and routing fees. This paper explores differences within pathfinding strategies used by prominent Lightning Network node implementations, which include different underlying cost functions and different constraints, as well as different greedy algorithms of shortest path-type. Surprisingly, we observe that the pathfinding problems that most LN node implementations attempt to solve are NP-complete, and cannot be guaranteed to be optimally solved by the variants of Dijkstra’s algorithm currently deployed in production. Through comparative analysis and simulations, we evaluate efficacy of different pathfinding strategies across metrics such as success rate, fees, path length, and timelock. Our experiments indicate that the strategies used by Eclair are advantageous in terms of payment reliability and result in paths with low fees. LND exhibits moderate success rates, while LDK results in paths with higher fee levels for smaller payment amounts; furthermore, CLN stands out for its minimal timelock paths. Additionally, we investigate the impact of Lightning node connectivity levels on routing efficiency. The findings of our analysis provide insights towards future improvements of pathfinding strategies and algorithms used within the Lightning Network.
@@ I Introduction and Background
By leveraging the trust and security of the Bitcoin blockchain in its payment channels, the Lightning Network is designed to offer bitcoin-denominated fast, cost-effective, and scalable transactions. Within the LN, users establish bidirectional payment channels, which collectively form a peer-to-peer network. These channels operate as 2-of-2 multisignature contracts between two parties (nodes) that hold bitcoin and negotiate ownership of those coins. Through these channels, users, which serve as nodes within the LN, can conduct off-chain payments. Settlement of these off-chain payments on the Bitcoin blockchain occurs only when necessary, such as when a user opts to close a channel. This architecture renders the Lightning Network scalable and enables instant payments [5].
@@
Multi-hop payments are facilitated through source routing, wherein the source locates a path to the receiver for payment routing. Initially, the sender receives a cryptographic hash from the recipient. For the sender’s payment to succeed, the recipient, upon receiving the HTLC, must return the pre-image along the HTLC-routed path within a specified timelock. Consequently, upon payment initiation, a series of HTLCs are established along the intended path. Each HTLC represents a commitment of funds from the preceding node to the next node, contingent upon the fulfillment of cryptographic proof. Once all contracts are established, the receiver shares the pre-image of the hash with its predecessor hop, which forwards it along the path until it reaches the sender, thereby resolving HTLCs along the path [8]. LN nodes can route payments only if they have sufficient balance on their side of the channel. A key practical challenge for successful payments within the LN is that the sender node selecting a payment path only possesses information about the capacity of channels, but not about the balance distribution within them [9, 10], which might lead to the selection of payment path that cannot successfully route the payment amount.
In order to discover paths for routing payments, the sender node utilizes a pathfinding algorithm. Popular and widely used pathfinding algorithms have been developed by various LN implementations, including Lightning Network Daemon (LND) [11], Core Lightning (CLN) [12], Lightning Development Kit (LDK) [13, 14], and Eclair [15]. While mostly using a variant of a shortest-path algorithm, each of these node implementations differs in its pathfinding strategy in substantial details such as in the weight function that establishes underlying path costs. With their different choices, these LN client implementations adopt diverse approaches to the pathfinding process with different trade-offs regarding the fees incurred by payments, the typical length of payment paths, incurred delays and payment reliability. Only little research has been dedicated to selecting the appropriate client (or pathfinding strategy) based on different need profiles regarding these trade-offs inherent to payment delivery within LN.
@@ II Related Work
The study by [10] concentrates on the reliability and privacy of LN payments by introducing a mathematical framework to model uncertainty in channel balances using probability theory. Furthermore, [16] explores the privacy properties of the Lightning Network through simulations. [17] investigates the structure of the Lightning Network and the routing algorithms used by LND, CLN and Eclair. Then it explores the effects of routing choices by analyzing the feasibility of Denial-of-Service attacks and route hijacking across these three implementations. In our work, we specifically analyze the effects of design elements in the pathfinding algorithm and evaluate their empirical performance from an operational perspective.
[18] examines the potential presence of the Braess paradox in off-chain networks and proposes solutions to mitigate its impact through analysis employing three routing algorithms, including LND and CLN. [19] analyzes the Lightning Network’s topology and introduces a strategy called ProfitPilot for connecting new nodes in payments by encouraging cycle creation in payment channel networks. Studies such as [20, 16, 21, 22] present topological analyses of the Lightning Network. In our study, we utilize insights into the Lightning Network topology, including the node connectivity levels, to analyze the performance of pathfinding strategies for different network participants.
@@
Beyond the scope of our exposition are payment delivery strategies based on (atomic) multi-part payments [6, Chapter 12], which have been introduced as an extension to the original LN protocol [26]. [27] argues that multipart payment delivery can benefit from a minimum cost flow formulation, which generalizes the concepts of payment paths, instead of following a shortest-path framework. [28] proposes a variation of Boomerang which requires less computing power and latency for faster multi-path payments. While multi-part payments are very promising, they do currently not (yet) represent the mainstream practice in all LN implementations, where single-path payments still dominate. Thus, our study focuses on single-path payments, but the conducted analysis could be extended to multi-part payments in future work.
@@ III Pathfinding and Route Selection
For the purpose of pathfinding, the Lightning Network can be modeled as a directed graph $G=(V,E)$, in which each network participant is represented as a node or vertex $v\in V$, and a payment channel between two nodes $u,v\in V$ is represented as the pair of edges $e=(u,v)\in E$ and $e=(v,u)\in E$. Each edge in the graph is associated with specific channel attributes such as fees, capacity, timelock delta, age, and information amount the maximal and minimal amount of currency units that can be settled in an HTLC. These channel attributes are also utilized as design elements for pathfinding purposes. It is important to note that the LN graph evolves dynamically over time–the number of nodes and channels changes through channel opening and closing transactions, and beyond that, node operators adapt channel parameters such as fee parameters and communicate those changes via a gossip protocol to other network participants.
@@
When the payment is initiated, the source node needs to select a path to the receiver node using pathfinding strategies on the LN graph. Existing LN clients employ instantiations of shortest path algorithms, such as Dijkstra’s algorithm, modified Dijkstra-style algorithms or Yen’s $K$-shortest path algorithm, to find the path between specified source and receiver nodes [30], [31], [32] and [33]. We refer to Section V for a discussion of these variants in the context of their use in LN implementations.
@@ IV Weight Function Design Elements
Given a sender node $s\in V$ and a recipient node $r\in V$, shortest path algorithms aim to find a path $p=(e_{1},\ldots,e_{\operatorname{length}(p)})$ of edges $e_{i}\in E$, where $s$ is the initial node of $e_{1}$ and $r$ is the terminal node of $e_{\operatorname{length}(p)}$, which minimizes a total path cost $c(p)$ given as the sum of edge weights given by a function $\operatorname{weight}:E\to\operatorname{\mathbb{R}}_{\geq 0}$ that maps to non-negative real numbers $\operatorname{\mathbb{R}}_{\geq 0}$, such that
@@
cf. [34, Chapter 22]. Recalling that an edge in the LN graph corresponds to a payment channel, this means that current LN client implementations, which use the shortest path algorithms to find their payment paths (see Section V), implicitly assume additive path costs in the sense of (1).
@@ IV-A Fees
When routing payments from a source $s$ to a recipient node $r$, the payment path $p$ typically involves intermediary nodes and multiple hops. These intermediary nodes charge fees for each successfully settled payment they participate in. These fees can be subdivided into two distinct parts: the base fee ($\mathrm{b}_{e}$), which remains constant regardless of the transaction size, and the proportional fee or fee rate ($\mathrm{f}_{e}$), which varies based on the amount being forwarded through the node. Consequently, the total fee charged by an intermediary node for forwarding a specific payment amount ($\mathrm{amt}_{e}$) is determined by the combination of these two components, such that
$\mathrm{fee}_{e}=\mathrm{b}_{e}+\mathrm{amt}_{e}\cdot\mathrm{f}_{e}.$ (3)
Since it is reasonable for the source node to try to find a path with fees as low as possible, one of the weight terms $a_{e}^{i}$ in (2) can be chosen as $a_{e}^{i}=\mathrm{fee}_{e}$ defined above. This choice is a primary design consideration in all LN implementations when formulating their weight functions. Since fees are directional, i.e., they can vary depending on the direction of payment flow between the same two nodes $u,v\in V$, a modeling of the LN as a directed graph instead of an undirected one is more appropriate. We also note that a purely edge-dependent fee as modeled in (3) is not fully realistic, as the payment amount $\mathrm{amt}$ needs to cover fees charged by routing nodes along the selected path, making $\mathrm{amt}$ path-dependent. Since the fees are typically several orders of magnitudes smaller than the amount to be received by the receiver node, we neglect this effect, in accordance with the practice in current LN node implementations.
@@ IV-B Timelock
Once a specific number of blocks is reached—referred to as a timelock—each incoming HTLC in a channel expires. This timelock serves as a time buffer within which the node aims to complete all the HTLC exchanges for a given transaction. For the length of the timelock, funds are locked in the payment channels. The total timelock of an HTLC, in direction from sender to recipient, depends on the timelock of the subsequent HTLC along a payment path plus the timelock difference value $\Delta^{\text{CLTV}}_{e}$, called timelock delta or CLTV expiry delta [^1] CLTV is an acronym of the Bitcoin script operator OP_CHECKLOCKTIMEVERIFY [6, Glossary]. that is set by the node $v$ at the outgoing side of the channel $e=(u,v)$ [6, Chapter 8]. In order to ensure that funds are locked for a shorter duration, it is reasonable for routing clients to prioritize the inclusion of channels with lower timelock deltas into the payment path compared to channels with higher timelock deltas.
@@ IV-C Success Probability
Due to the fact that payment paths in which the desired payment amount can be settled reliably are preferred, it is common to take an estimate of a success probability, which estimates the likelihood of a successful payment through a payment path, into account when searching for suitable payment paths $p$. In this subsection, we first discuss two different approaches for incorporating payment path success probabilities into a model for a total path cost, and then shed light on different approaches for modeling individual channel success probabilities.
@@ IV-C1 Path Success Probability
In existing literature on LN pathfinding, there are two different approaches for incorporating a path success probability estimate $P_{p}$ into a total path cost $c(p)$ of a payment path $p$—one with an additive term inversely proportional to $P_{p}$ and the second proportional to the negative logarithm of $P_{p}$.
The first, which is used in LND [35, 36] implies a success probability penalization that is proportional to the inverse $P_{p}^{-1}$ of the path success probability $P_{p}$. It is based on the following lemma, the proof of which is provided in [36].
@@
We note that in [35, 36], the above lemma is derived assuming the existence of a virtual cost $c_{\text{attempt}}$ only in case of a payment failure, but both assumptions lead to the same objective (4). In (4), the fee cost term $F_{p}$ fits well into the additive framework of (1) due the fact that routing fees accumulate additively along the path such that $F_{p}=\sum_{e\in p}\operatorname{weight_{a}}(e)$, where $\operatorname{weight_{a}}:E\to\operatorname{\mathbb{R}}_{\geq 0}$ is a non-negative, additive weight function.
However, for second summand of (4), the situation is different as an additive decomposition of $\frac{c_{\text{attempt}}}{P_{p}}$ across edges can be mathematically not justified. On the other hand, a reasonable assumption is that a payment across a path $p$ is successful if and only if the routing across all channels $e\in p$ is successful, who themselves are assigned channel success probabilities $P_{e}$ [10]. If we assume that the routing success across one channel $e\in p$ is independent from the routing success across any other channels along each path, we have that
$\begin{split}P_{p}&=P\left(\{\text{payment across path }p\text{ is successful}\}\right)\\ &=P\left(\cap_{e\in p}\{\text{payment across channel }e\text{ is successful}\}\right)\\ &=\prod_{e\in p}P_{e},\end{split}$ (5)
@@
Unfortunately, it is not possible to use a standard shortest path algorithm such as Dijkstra’s algorithm to find the path with the lowest additive-plus-multiplicative cost (6) as these algorithms all assume additive cost as in (1) [34, Chapter 22]. The LND LN node implementation mitigates this issue by modifying the priority queue used in Dijkstra’s algorithm to be adapted to (6). However, as we discuss in Section V, it can be shown that such a modified Dijkstra is not guaranteed to find the optimal path with respect to (6).
@@
which leads to a fully additive cost of the type (1) motivates the choice of a weight term $a_{e}^{i}=-\log(P_{e})$ in (2). Since the edge success probabilities $P_{e}$ are between 00 and $1$, these weight terms are always non-negative. If used without any other weight terms (e.g., fee-related ones), the resulting problem type corresponds to finding a maximally reliable path and is known to be equivalent to a shortest path problem [37, Exercise 4.39]. CLN, LDK and Eclair use this logarithmic success probability term in their weight function to prioritize paths with higher success probability within their pathfinding algorithms.